Systems Staff Meeting Notes 08/11/03

Systems Staff Meeting Notes
August 11, 2003

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart
Announcements
  1. Ailing critter discussion.
  2. New Xerox machine – no banding on copies.
  3. Hot locks on doors in room 450 and 3rd floor staff lounge. No explanation as to the cause.
  4. Faceplate is missing from the plug outlet in the Systems conference room. No explanation as to why.
  5. Sunshine yesterday!
  6. Suzy is out on jury duty. No word on how long.
Review Minutes of 08/04/03
  1. No word from Dell about the missing CPU from the West loading dock. Geoffrey has another call into Dell. Bill told Bill Hanssen to save the video from the delivery.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Nothing major on order.
  2. Nothing major broken.
Liaisons
  1. Scan spot checks: about half are fixed. One area liaison swore that he had done them all – checked four machines, three were undone. One had 17 critical updates to install. Found one where the machine thought the update was done, but the scan did not. Rebooting cleared it. As of Friday, there were 200-300 workstations still not up to date on patches.
  2. There has been some misconfiguration with IP addresses. Private IPs were using regular IPs. The upcoming move by West will require that all possible IPs are available. We are getting a decent response from the liaisons in cleaning up.
Networking
  1. Network support for renovation – no VOIP in Documents. We are now making adjustments to this. Bobbie and Tim went over and looked at the area, talked to an electrician. The ceiling grid cannot be installed until September 1st. The move is scheduled to take place on October1st – Net Services is checking on their availability, but they are busy. The total number of workstations moving from West to MSL is not known. Can’t estimate the public area until the numbers are known. Bill has not received any answers.
  2. Book Detention Facility – Waldo Rd – unknown what will be going out there or what type of networking will be necessary. Bill will meet tomorrow with Rob Roberts (coordinating West move) for more information. Hang and Rob estimate 12 people moving out there. Circulation functionality is a questionable issue.
  3. Tower, fourth floor East– need to run fiber, old copper may not work. Can run fiber up from room 16, but we’ll need to know how fussy they’re going to be about wires, noise, etc. No resolution on closet location. Tim will do a general estimate to get fiber to room 450 – where Personnel is going. New switches may need to be purchased. Port survey needed for East north tower. 92 people will move from West to East. That means 92 IPs and ports. VLAN at the core to hook 238, 239 & 223. Technical reasons as to why it may or may not be possible to move the IPs. Facilities said they are moving as soon as possible, but not all at once. Phones will be another problem – our local building switch will not be able to accommodate all of the people. Barbara Oliver will be disappearing until Labor Day – she is the financial manager for the renovation.
  4. Service Pack 4 – rolling out nicely on public machines. Some people are having difficulty reading and understanding the instructions that we’ve posted. Geoffrey and Marvin have been working on West, Paul and Leilani are helping in Music and Education. It is a long process – can take up to 2 hours per machine. Many machines had not been updated since November 2002. File and print sharing is documented in the instructions.
Software Projects Status
  1. SysRequest – current version is ready to add attachments – has to go on NT8. It is currently sending email notification to users and approvers when the status is changed or updated.
  2. FilmLog – all requests are done, it is in the testing phase. Docutek licensing: sent them Bill’s suggestions – they are reviewing and will advise.
  3. Book/Film Replacement is moving slowly.
  4. ILLIAD – still tweaking, it keeps needing maintenance.
DP Coordinator
  1. It is currently the 2-week break before the fall semester. Nikia will be working.
  2. Fall student line-up will be: Nikia, Laura & Nicole.


Back to the Systems Homepage

Last updated August 18, 2003
by Debra Fetzer
debharr@mail.uflib.ufl.edu